OpenStudy (anonymous):

Your cap size is based on the circumference of your head (in inches). For a certain range of head circumferences c, cap size s can be modeled by the following equation: S= c-1/3 If you wear a size 7.25 cap, what is your head circumference?

OpenStudy (anonymous):

7.58

OpenStudy (anonymous):

how did you come up with 7.58

OpenStudy (anonymous):

c = S + 1/3 c = 7.25 + 0.33 = 7.58

OpenStudy (anonymous):

welcome missy

OpenStudy (anonymous):

ok i c how you did it now thanks a lot